# Phaedrus: Predicting Dynamic Application Behavior with Lightweight Generative Models and LLMs

### v3

@@ 4 SyncCacheViewAutPix(); // cold
Listing 3: Simplified Code snippet from imagick in SPEC2017 (Bucek et al., 2018). The call-graph traversal from hot kernels allows Dynamis to isolate ‘moderately’ hot functions
@@ 5. Evaluation
Experimental Setup: The experiments were conducted on a Dell 7820 Precision Tower with Intel Xeon Gold 6230R processors. The Intel system has 2 sockets (13 cores each), 36 MB L3-cache, running Ubuntu 20.04.6 LTS. The compiler passes and algorithms in front-end of Morpheus & Dynamis were implemented in LLVM 15.0.0, while Pytorch version 2.4.0, was used in Morpheus’s deep-learning back-end with Python 3.8.10. GPT-5, Claude-4-Sonnet and GPT 4.0 were used as the LLM in Dynamis.
@@ 5.3. Leveraging Phaedrus: Profile-Guided Optimization
To investigate this effect further, we analyzed the contents of the predicted profdata and the ground truth closely. Profiling tool gprof highlights expensive compute-bound routines that dominate wall-clock time, while profdata emphasizes fast but frequently executed functions. It became evident that functions dominating runtime are not always those most amenable to direct optimization. In some cases, performance can be improved indirectly—by transforming other functions that invoke these heavy kernels. For instance, in 520.omnetpp_r520.omnetpp\_r, the low-level function cMessageHeap::shiftup predicted by Dyanmis appeared trivial in gprof, yet it dominated execution counts in profdata. We observed that Dynamis’ reasoning consistently balanced between capturing both computationally intensive and optimization-relevant routines. In spite of being prompted to identify runtime-dominated hotspots, Dynamis still surfaced functions that were most likely to yield measurable performance improvement when optimized. This effect was also observed in the application binary sizes, where Dyanmis achieved an improvement of 5.19 over O3. This is achieved by transforming the function layout within the application binary. This optimization reduces the binary size by enabling superior function compression (Hoag et al., 2024), and can potentially improve performance as the new layout would minimize page faults by appropriately ordering them.
@@ 5.4. Case Study: Real-World Benchmarks
In order to better demonstrate the capabilities of Dynamis on larger real world programs, we evaluated the framework on 8 popular graph processing problems ((Beamer et al., 2017)) (BFS, CC, CC_SV, TC, PR, PR_SPMV, SSSP) with an undirected graph of degree 15 with 16777216 nodes and 268435177 undirected edges, and also with the popular compiler benchmark GCC, which has 1304 KLOC. Fig. 19 shows that Dynamis outperforms O3 baseline by 4.45% for GCC. It navigated the huge codebase by starting with the program inputs & flags, based on which it inferred the dominant compiler passes. Then it identified the ”known” heavy kernels within those passes (bitmaps and graph operations, points-to analysis, etc), which were then ranked by their expected algorithmic cost. This shows how Dynamis is able to adapt its reasoning for large codebase by tracing its thinking from the inputs.
@@ 5.5. Evaluating Morpheus: Application Profile Generalization
$\star$ Need for Compressing WPP based Function Profiles: Table 21 shows the total tokens and disk size for WPP based function profiles expressed in terms of application function calls for SPEC2k17 (Bucek et al., 2018) benchmarks. As shown, the total function call tokens can range up to $10^{10}$ tokens per profile while incurring disk sizes of a few GB to $10^{2}$ GB. Without the compression techniques employed by Morpheus, such profiles cannot be analyzed. For instance, Fig. 23 summarizes the result of training a simple RNN model with one hidden layer on the uncompressed WPP data generated from the smallest program input, in SPEC2k17 Benchmarks. Out of 14 benchmarks, the training process on 7 benchmarks failed because of exceeding the tensor’s default memory capacity in the Pytorch framework, while 5 benchmarks had out-of-memory errors, and were terminated by the OS scheduler. Furthermore, Fig. 23 depicts the low ratio of unique tokens in WPP based function profiles. This indicates a significantly high rate of duplicate (and redundant) tokens, which can lead to excessive bias and unnecessary computations during training.
@@ 6. Related works
Recent research in code optimization has transitioned from rule-based compiler heuristics to large language model (LLM)–driven code transformation and refactoring (Gao et al., 2024; Rosas et al., 2024; Ma et al., 2024a; Pomian et al., 2024; Cordeiro et al., 2024). LLM frameworks combine sequence generation with iterative refinement using evolutionary search, adaptive retrieval, and specialized prompting (Liu et al., 2023a; Wu et al., 2024). Comparative studies show that LLMs can rival or surpass traditional compilers such as CETUS and PLUTO in optimization quality but often struggle with correctness and scalability on large real-world programs. Instruction-tuned models (Ma et al., 2024a) further address these issues through optimization-aware fine-tuning, improving reliability and consistency across diverse workloads.
@@
In parallel, machine learning–based static profiling has emerged as an alternative to dynamic PGO, eliminating the need for costly runtime instrumentation and representative input collection. VESPA (Moreira et al., 2021) pioneered static frequency prediction for binary optimization in Meta’s BOLT (Panchenko et al., 2018), achieving up to 6% speedup over Clang –O3 by estimating basic-block frequencies directly from static features. GraaSP (Čugurović and Janičić, 2024) extended this concept to compiler integration within GraalVM (Šipek et al., 2020), embedding ML inference during compilation to remove the two-phase PGO cycle and delivering a 7.5% runtime gain. GraalNN (Milikic et al., 2025) advanced the field further by employing Graph Neural Networks to model call-graph context, refining predictions across parsing and inlining phases for greater accuracy. Building on these insights, our work explores input-specific static profiling without execution, using LLMs to infer hot regions based on algorithmic structure, input size, and termination behavior. Unlike prior systems targeting speed in Java or binary infrastructures, our framework focuses on code optimization in C/C++ workloads, evaluated against PGO baselines on SPEC CPU 2017. Though not directly comparable, our results demonstrate that LLMs can approximate—and often surpass—traditional profiling accuracy while offering fine-grained, input-aware guidance.
@@
On the other hand, the concept of capturing comprehensive dynamic program behavior through Whole Program Paths (WPP) has been extensively explored in the compiler literature (Larus, 1999; Zhang and Gupta, 2001; Burtscher and Jeeradit, 2003; Zhang and Gupta, 2002; Law and Rothermel, 2003; Milenkovic and Milenkovic, 2003; Zhang and Gupta, 2004; Lin et al., 2007). WPPs provide a single, compact description of a program’s entire dynamic control flow, including interprocedural calls and loop iteration, which is essential for finding ”hot subpaths” for optimization. While techniques for compressing WPP profiles are well-developed, they are inherently limited to analyzing a single profile instance generated from one specific input data set. Our approach, in contrast, attempts to generalize across multiple profiles by using a novel, lossless compression scheme that is consistent across different inputs, enabling the prediction of program behavior for unseen execution scenarios.

### v4

@@ 4 SyncCacheViewAutPix(); // cold
Listing 3: Simplified Code snippet from imagick in SPEC2017 (bucek2018spec). The call-graph traversal from hot kernels allows Dynamis to isolate ‘moderately’ hot functions
@@ 5. Evaluation
Experimental Setup: The experiments were conducted on a Dell 7820 Precision Tower with Intel Xeon Gold 6230R processors. The Intel system has 2 sockets (13 cores each), 36 MB L3-cache, running Ubuntu 20.04.6 LTS. The compiler passes and algorithms in front-end of Morpheus & Dynamis were implemented in LLVM 15.0.0, while Pytorch version 2.4.0, was used in Morpheus’s deep-learning back-end with Python 3.8.10. GPT-5, Claude-4-Sonnet and GPT 4.0 were used as the LLM in Dynamis. The LLMs were invoked in an interactive environment by Cursor (cursor_ai_2024) .
@@ 5.3. Leveraging Phaedrus: Profile-Guided Optimization
To investigate this effect further, we analyzed the contents of the predicted profdata and the ground truth closely. Profiling tool gprof highlights expensive compute-bound routines that dominate wall-clock time, while profdata emphasizes fast but frequently executed functions. It became evident that functions dominating runtime are not always those most amenable to direct optimization. In some cases, performance can be improved indirectly—by transforming other functions that invoke these heavy kernels. For instance, in 520.omnetpp_r520.omnetpp\_r, the low-level function cMessageHeap::shiftup predicted by Dyanmis appeared trivial in gprof, yet it dominated execution counts in profdata. We observed that Dynamis’ reasoning consistently balanced between capturing both computationally intensive and optimization-relevant routines. In spite of being prompted to identify runtime-dominated hotspots, Dynamis still surfaced functions that were most likely to yield measurable performance improvement when optimized. This effect was also observed in the application binary sizes, where Dyanmis achieved an improvement of 5.19 over O3. This is achieved by transforming the function layout within the application binary. This optimization reduces the binary size by enabling superior function compression (hoag2024reordering), and can potentially improve performance as the new layout would minimize page faults by appropriately ordering them.
@@ 5.4. Case Study: Real-World Benchmarks
In order to better demonstrate the capabilities of Dynamis on larger real world programs, we evaluated the framework on 8 popular graph processing problems ((beamer2017gapbenchmarksuite)) (BFS, CC, CC_SV, TC, PR, PR_SPMV, SSSP) with an undirected graph of degree 15 with 16777216 nodes and 268435177 undirected edges, and also with the popular compiler benchmark GCC, which has 1304 KLOC. Fig. 19 shows that Dynamis outperforms O3 baseline by 4.45% for GCC. It navigated the huge codebase by starting with the program inputs & flags, based on which it inferred the dominant compiler passes. Then it identified the ”known” heavy kernels within those passes (bitmaps and graph operations, points-to analysis, etc), which were then ranked by their expected algorithmic cost. This shows how Dynamis is able to adapt its reasoning for large codebase by tracing its thinking from the inputs.
@@ 5.5. Evaluating Morpheus: Application Profile Generalization
$\star$ Need for Compressing WPP based Function Profiles: Table 21 shows the total tokens and disk size for WPP based function profiles expressed in terms of application function calls for SPEC2k17 (bucek2018spec) benchmarks. As shown, the total function call tokens can range up to $10^{10}$ tokens per profile while incurring disk sizes of a few GB to $10^{2}$ GB. Without the compression techniques employed by Morpheus, such profiles cannot be analyzed. For instance, Fig. 23 summarizes the result of training a simple RNN model with one hidden layer on the uncompressed WPP data generated from the smallest program input, in SPEC2k17 Benchmarks. Out of 14 benchmarks, the training process on 7 benchmarks failed because of exceeding the tensor’s default memory capacity in the Pytorch framework, while 5 benchmarks had out-of-memory errors, and were terminated by the OS scheduler. Furthermore, Fig. 23 depicts the low ratio of unique tokens in WPP based function profiles. This indicates a significantly high rate of duplicate (and redundant) tokens, which can lead to excessive bias and unnecessary computations during training.
@@ 6. Related works
Recent research in code optimization has transitioned from rule-based compiler heuristics to large language model (LLM)–driven code transformation and refactoring (gao2024searchbasedllmscodeoptimization; rosas2024aioptimizecodecomparative; ma2024llamocoinstructiontuninglarge; pomian2024furtherllmsidestatic; cordeiro2024empiricalstudycoderefactoring). LLM frameworks combine sequence generation with iterative refinement using evolutionary search, adaptive retrieval, and specialized prompting (liu2023promptingframeworkslargelanguage; wu2024evolutionarycomputationeralarge). Comparative studies show that LLMs can rival or surpass traditional compilers such as CETUS and PLUTO in optimization quality but often struggle with correctness and scalability on large real-world programs. Instruction-tuned models (ma2024llamocoinstructiontuninglarge) further address these issues through optimization-aware fine-tuning, improving reliability and consistency across diverse workloads.
@@
In parallel, machine learning–based static profiling has emerged as an alternative to dynamic PGO, eliminating the need for costly runtime instrumentation and representative input collection. VESPA (vespa2021moreira) pioneered static frequency prediction for binary optimization in Meta’s BOLT (panchenko2018boltpracticalbinaryoptimizer), achieving up to 6% speedup over Clang –O3 by estimating basic-block frequencies directly from static features. GraaSP (vcugurovic2024graalsp) extended this concept to compiler integration within GraalVM (graalvm2020), embedding ML inference during compilation to remove the two-phase PGO cycle and delivering a 7.5% runtime gain. GraalNN (milikic2025graalnn) advanced the field further by employing Graph Neural Networks to model call-graph context, refining predictions across parsing and inlining phases for greater accuracy. Building on these insights, our work explores input-specific static profiling without execution, using LLMs to infer hot regions based on algorithmic structure, input size, and termination behavior. Unlike prior systems targeting speed in Java or binary infrastructures, our framework focuses on code optimization in C/C++ workloads, evaluated against PGO baselines on SPEC CPU 2017. Though not directly comparable, our results demonstrate that LLMs can approximate—and often surpass—traditional profiling accuracy while offering fine-grained, input-aware guidance.
@@
On the other hand, the concept of capturing comprehensive dynamic program behavior through Whole Program Paths (WPP) has been extensively explored in the compiler literature (larus1999whole; zhang2001timestamped; burtscher2003compressing; zhang2002path; law2003whole; milenkovic2003stream; zhang2004whole; lin2007design). WPPs provide a single, compact description of a program’s entire dynamic control flow, including interprocedural calls and loop iteration, which is essential for finding ”hot subpaths” for optimization. While techniques for compressing WPP profiles are well-developed, they are inherently limited to analyzing a single profile instance generated from one specific input data set. Our approach, in contrast, attempts to generalize across multiple profiles by using a novel, lossless compression scheme that is consistent across different inputs, enabling the prediction of program behavior for unseen execution scenarios.